OpenStudy (countrygirl1431):

Write the standard form of the equation of the line that passes through the given points. Include your work in your final answer. Type your answer in the box provided or use the upload option to submit your solution. (5, 3) and (8, -2)

OpenStudy (countrygirl1431):

@sweetburger

OpenStudy (sweetburger):

Ok our first step will be to find the slope because 2 sets of points are given

OpenStudy (sweetburger):

So our slope is 

OpenStudy (countrygirl1431):

ok i got that much

OpenStudy (sweetburger):

The next step is to use point slope form which is y - y1 = m(x-x1) and we substitute m= -5/3 into the equation and we use the set of points (5,3) (although it doesnt matter which set of points you pick) and we plug those values into the equation. This ends up looking like this y-3=(-5/3)(x-5)

OpenStudy (texaschic101):

or you can just use y = mx + b form

OpenStudy (sweetburger):

very true

OpenStudy (texaschic101):

either way will get you to the same answer

OpenStudy (texaschic101):

sorry...didn't mean to interrupt....you all can continue now...I will be quiet

OpenStudy (sweetburger):

no worries @texaschic101

OpenStudy (sweetburger):

So anyways we must condense y-3=(-5/3)(x-5) so that looks like this